# Model Tensor Planning

### v1

@@ 4 Experiments
  • Further ablations on sweeping graph parameters $M,N$, softmax weighting, and JAX performance benchmark are presented in Section A.5.
We study the cumulative cost $J({\bm{\tau}},{\bm{\mu}})$ Eq. 2 over the control timestep for each task. For each experiment, we take the minimum cumulative cost in batch rollouts at each timestep, and plot the mean and standard deviation over $5$ seeds.
Practical Settings. All algorithms and environments are implemented in MuJoCo XLA (Todorov et al., 2012; Kurtz, 2024), to utilize the jit and vmap JAX operators for efficient vectorized sampling and rollout on multiple model instances. All experiment runs are sim-to-sim evaluated (MuJoCo XLA to MuJoCo). In particular, we introduce some modeling errors in MuJoCo XLA. Then, for online domain randomization, we randomize a set of $R$ models $\{f_{j}({\bm{x}},{\bm{u}})\}_{j=1}^{R}$, then we perform batch sampling and rollout of $R\times B$ trajectories. Finally, the cost evaluation is averaged on the $R$ domain randomization dimension. Note that, for this paper, we deliberately design the task costs to be simple and set sufficiently short planning horizons to benchmark the exploratory capacity of algorithms. In practice, one may design dense guiding costs to achieve the tasks. Further task details are presented in Section A.3.
@@ 4.1 Motivating Example
Here, we provide an experimental analysis of the baselines’ exploration capacity on Navigation environment (cf. Fig. 3), where the point-mass agent is controlled by an axis-aligned $2$-dim velocity controller. We compare MTP-Bspline and MTP-Akima to evolutionary algorithms and standard MPC baselines, with maximum sampling noise settings (cf. Fig. 3). In particular, given the control limits $[-1,1]$ on x-y axes, we set the standard deviation $\sigma=1$ for sampling noise of MPPI and PS, and population generation noise for OpenAI-ES and DE.
Figure 3: Motivation comparison of MTP methods versus baselines with $B=256$ on Navigation environment. The environment is designed to be challenging to reach the green goal, requiring strong exploration to avoid large local minima in the middle (see task details in Section A.3). We plan with $T=20$ with $\Delta t=0.05s$. The figures show $5$ random traces of white rollouts. (Left) MTP-Akima rollouts reach the green goal very early due to high-entropy tensor sampling, while (Right) OpenAI-ES struggles to generate a rollout exploring the way out of large local minima.
Fig. 3 also shows the cost convergence and the cost entropy curves over timesteps, in which the entropy is computed as $H=-\sum_{j=1}^{B}P_{j}\log P_{j},\,P_{j}=\exp(J_{j})/(\sum_{l}\exp(J_{l}))$, where $\{J_{j}\}_{j=1}^{B}$ is the batch of cumulative rollout costs. The entropy represents the diversity of rollout evaluation, implying the exploration capability of algorithms. We observe that MTP-Akima has the highest entropy curve, inducing the lowest cost convergence over timesteps, while other baselines converge to the middle minima. In this case, MTP-Bspline also struggles due to conservative interpolation in tensor sampling, achieving moderate entropy, yet higher than evolutionary strategies.
@@ 4.2 Comparison Experiments
We analyze the performance comparison of MTP and the baselines over a wide range of robotics tasks representing different dynamics and planning cost settings (cf. Fig. 4). In each task, we tune the baselines and set the same white noise standard deviation for MTP/MPPI/PS to study the performance gain by tensor sampling, while using the default hyperparameters for evolutionary baselines.
Comparison Environments. PushT (Chi et al., 2023), Cube-In-Hand (Andrychowicz et al., 2020) require intricate manipulation environments where robust exploration is critical due to complex contact dynamics and precise multi-step manipulation requirements. G1-Standup, G1-Walk represent high-dimensional robotic tasks demanding substantial computational resources and sophisticated control strategies. Crane, Walker (Towers et al., 2024) present underactuated and nonlinear dynamic challenges. To ensure fair comparison, for all baselines, we fix the same number of rollouts $B=16$ on Crane, and $B=128$ for all other tasks. All tasks are implemented in hydrax (Kurtz, 2024). Further experiment details are in Section A.4.
Figure 4: Performance comparison of MTP variants against standard MPC methods (MPPI, PS) and evolutionary algorithms (OpenAI-ES, DE). The gray dashed line depicts the task success bar, while the red represents the timestep such that the first algorithm statistically succeeds the task, or fails last in G1-Walk.
@@
In Crane environment, we apply heavy modeling errors of mass, inertia, and pulley/joint damping. MTP variants excel by maintaining stable control trajectories with smooth transitions, effectively navigating the nonlinear dynamics and underactuation. The B-spline interpolation’s conservative nature helps avoid overshooting and instability prevalent in these tasks, thus outperforming both evolutionary algorithms and standard MPC methods that tend to produce erratic control inputs. In Walker, we apply no modeling error, and thus the performance of MTP variants and MPPI, PS are similar.
@@ 4.3 Design Ablation
We conduct an ablation study analyzing the effect of varying two crucial hyperparameters—the number of elites $E$ and the mixing rate $\beta$—on MTP algorithmic performance. Fig. 5 presents heatmaps indicating accumulated cost over time for each task and interpolation method (MTP-Linear, MTP-Bspline, MTP-Akima). Each heatmap illustrates distinct algorithmic realizations at its corners. Specifically, the bottom-left corner represents PS, characterized by a single elite and purely white noise sampling. Conversely, the bottom-right corner corresponds to Predictive Tensor Sampling (TS-PS), maintaining a single elite but employing full tensor sampling. Due to the softmax update (Algorithm 2 Line 11-12), the top-left corner realizes MPPI (with adaptive sampling covariance), leveraging all candidate samples with local noise sampling, while the top-right corner reflects Tensor Sampling-MPPI (TS-MPPI), utilizing all candidates and full tensor sampling for maximum exploration.
Figure 5: Mixing rate $\beta$ and number of elites $E$ sweep on PushT, G1-Walk tasks with $B=128$ to investigate the algorithmic update rule Algorithm 2 Line 9-12. The heatmap indicates accumulated cost over timesteps at termination, and the heat value range is fixed for each task/row.
We observe the consistent pattern that MTP performance degrades at the extremes of $\beta$. In PushT, moderate values of $\beta$ lead to significantly lower costs, while the absence of tensor sampling ($\beta=0$) yields poor performance due to inadequate exploration. This observation reinforces the effectiveness of the $\beta$-mixing strategy for balancing global and local sampling contributions. Interestingly, the number of elites $E$ has a limited impact in PushT, likely due to the task’s insensitivity to control stability. However, in G1-Walk, the choice of $E$ is crucial. Using a single elite ($E=1$) , corresponding to the PS control scheme, leads to unstable and jerky behavior, which aligns with the poor performance observed in Fig. 4. At the other extreme, with $\beta=1$ (full tensor sampling), performance also degrades. This is attributed to the fixed rollout budget $B$; as $M$ increases, global samples become too sparse to effectively capture the fine-grained control required for stable gait tracking. In this case, exploitation with local samples is essential to maintain intricate motion tracking control.
@@ 4.4 Sensitivity Ablation
Here, we perform sensitivity analyses for various critical algorithmic hyperparameters. Fig. 6 evaluates sensitivity to the mixing rate $\beta$ for different tasks. For the PushT environment, results show minimal sensitivity across mixing rates, as the task inherently lacks significant failure modes, ensuring consistent success regardless of the exploration-exploitation balance. In contrast, the Cube-In-Hand task demonstrates high sensitivity, with larger mixing rates causing instability due to the cube falling out of grasp frequently. Optimal performance is thus achieved with lower $\beta$ values, suggesting careful management of exploration intensity. Furthermore, for the high-dimensional G1-Standup task, a smaller mixing rate helps stabilize the control, enabling the robot to achieve a more consistent and stable stand-up performance.
@@ 5 Related Works
We review related efforts across two major directions: the vectorization of sampling-based MPC and sampling-based motion planning. While these approaches originate from different planning paradigms—dynamics-aware MPC versus collision-free geometric planning—they share a common structure: interacting with an agent-environment model (e.g., dynamics model or collision checker) that can be vectorized for efficient, batched computation. Both domains benefit from high-throughput sampling, making them increasingly amenable to modern GPUs/TPUs.
@@
Sampling-based MPC Vectorization. Sampling-based MPC (Mayne, 2014) has been successfully applied to high-dimensional, contact-rich control problems. Methods such as Predictive Sampling (PS) (Howell et al., 2022), Model Predictive Path Integral (MPPI) (Williams et al., 2017), and CEM-based MPC (Pinneri et al., 2021) rely on parallel sampling of control trajectories and subsequent rollouts using a system dynamics model. These methods naturally benefit from vectorized simulation backends, and recent works have extended them toward more structured and efficient exploration. For instance, inspired by the diffusion process, DIAL-MPC (Xue et al., 2024) enhances exploration coverage and local refinement simultaneously, achieving high-precision quadruped locomotion and outperforming reinforcement learning policies (Schulman et al., 2017) in climbing tasks. STORM (Bhardwaj et al., 2022) demonstrates GPU-accelerated joint-space MPC for robotic manipulators, achieving real-time performance while handling task-space and joint-space constraints. Other recent efforts integrate GPU-parallelizable simulators, such as IsaacGym (Makoviychuk et al., 2021), into the MPC loop (Pezzato et al., 2025) for online domain randomization, removing the need for explicit modeling and enabling real-time contact-rich control. In another line, CoVO-MPC (Yi et al., 2024) improves convergence speed by optimizing the covariance matrix in sampling, resulting in performance gains in both simulated and real-world quadrotor tasks. These advances demonstrate that structured, parallel control sampling can be effectively deployed in high-stakes robotics applications using vectorized dynamic models.
@@
Motion Planning Vectorization. Recent advances in sampling-based motion planning have shown that classical methods such as RRT (Kuffner & LaValle, 2000) can be significantly accelerated using parallel computation, while preserving theoretical guarantees like probabilistic completeness, which is another form of maximum exploration. Early work focused on accelerating specific subroutines like collision checking (Bialkowski et al., 2011; Pan & Manocha, 2012), but more recent efforts have restructured planners for full GPU-native execution. Examples include GMT* (Lawson et al., 2020), VAMP (Thomason et al., 2024), pRRTC (Huang et al., 2025), and Kino-PAX (Perrault et al., 2025), which achieve millisecond-scale planning in high-dimensional configuration spaces by parallelizing sampling, forward kinematics, and tree expansions. GTMP (Le et al., 2024) pushes this even further by implementing the sampling, graph-building, and search pipeline as tensor operations over batch-planning instances, showcasing the feasibility of real-time planning across multiple environments.
@@ 6 Discussions and Conclusions
In this work, we introduced Model Tensor Planning (MTP), a robust sampling-based MPC approach designed to achieve global exploration via maximum entropy sampling. Theoretically, we demonstrated that in the limits of infinite layers $M$ and samples per layer $N$, our tensor sampling method attains maximum entropy, thereby efficiently approximating the full trajectory space. Furthermore, MTP is intentionally designed to be practically feasible, enabling straightforward implementation for sampling high-entropy control trajectories (see Section A.5).
@@
Limitations.
While MTP demonstrates strong performance across diverse control tasks, it inherits several limitations typical of sampling-based methods. First, its computational cost scales with the number of rollouts, making it challenging to deploy on hardware with limited parallel computing. Second, while our tensor-based sampler improves exploration coverage, it does not leverage task-specific priors or learning-based proposal distributions, which could further improve sample efficiency. Finally, MTP relies on a fixed dynamics model and does not currently support model uncertainty, limiting its robustness in partially observed or stochastic environments.
Acknowledgments
This work was funded by the German Federal Ministry of Education and Research Software Campus project ROBOSTRUCT (01S23067).
@@ A.1 Proof of Theorem 1
Let $u\in{\mathbb{F}}$ be any path that is uniformly continuous and has bounded variation $TV(u)<\infty$. We begin by constructing a piecewise linear path approximating $u$, $g_{M}:[0,1]\to{\mathbb{U}}$ by dividing the interval $[0,1]$ into $M-1$ subintervals, i.e., $[t_{1},t_{2}],\ldots,[t_{M-1},t_{M}]$ with $10=t_{1}<t_{2}<\ldots<t_{M}=$ 1. On each subinterval $[t_{i},t_{i+1}]$, we define the corresponding segment of $g$ to approximate $u$
@@
Since $u$ is uniformly continuous on $[0,1]$, for any $\epsilon>0$, there exists $\delta>0$ such that for all $t,s\in[0,1]$, if $|t-s|<\delta$, then
@@
Let $u\in{\mathbb{F}}$ be any control path and ${\mathcal{G}}(M,N)$ be a random multipartite graph with $M$ layers and $N$ uniform samples per layer (cf. Definition 1). Assuming a time sequence (i.e., knots) $10=t_{1}<t_{2}<\ldots<t_{M}=$ 1 with equal intervals, associating with layers $L_{1},\ldots,L_{M}\in{\mathcal{G}}(M,N)$ respectively, then
@@
First, Lemma 2 implies that there exists a sequence of linear piecewise $g_{M}$, having $M-1$ equal intervals approximating $u$, converging to $u$ as $M\to\infty$.
Let $\hat{g}_{M}\in{\mathcal{G}}(M,N)$ be a control path in ${\mathcal{G}}$ (cf. Definition 4). Since the time sequence $10=t_{1}<t_{2}<\ldots<t_{M}=$ 1 corresponding to layers $L_{1},\ldots,L_{M}\in{\mathcal{G}}(M,N)$ has equal intervals, we can consider $\hat{g}_{M}$ having $M-1$ segments approximating $g_{M}$ without loss of generality.
@@
for some $c>0$. From Lemma 1, the probability that every layer contains at least one such sample, such that $\|g_{M}-\hat{g}_{M}\|_{\infty}<\epsilon$, is at least $1-Me^{-cN}$, which converges to 1 as $N\to\infty$.
From Lemma 2, for sufficiently large $M$, we have $\|u-g_{M}\|_{\infty}<{\epsilon}$. Now, due to ${\mathbb{U}}$ is compact, we can apply the triangle inequality as $N\to\infty$

### v2

@@ 4 Experiments
We study the cumulative cost $J({\bm{\tau}},{\bm{U}})$ Eq.˜2 over the control timestep for each task. For each experiment, we take the minimum cumulative cost in batch rollouts at each timestep, and plot the mean and standard deviation over 555 seeds. Further ablations on sweeping graph parameters $M,N$, softmax weighting, and JAX planning performance benchmark are presented in Section˜A.5.
Practical Settings. All algorithms and environments are implemented in MuJoCo XLA (Todorov et al., 2012; Kurtz, 2024), to utilize the jit and vmap JAX operators for efficient vectorized sampling and rollout on multiple model instances. All experiment runs are sim-to-sim evaluated (MuJoCo XLA to MuJoCo). In particular, we introduce some modeling errors in MuJoCo XLA. Then, for online domain randomization, we randomize a set of $R$ models $\{f_{j}({\bm{x}},{\bm{u}})\}_{j=1}^{R}$, then we perform batch sampling and rollout of $R\times B$ trajectories. Finally, the cost evaluation is averaged on the $R$ domain randomization dimension. Note that, for this paper, we deliberately design the task costs to be simple and set sufficiently short planning horizons to benchmark the exploratory capacity of algorithms. In practice, one may design dense guiding costs to achieve the tasks. Further task details are presented in Section˜A.3.
@@ 4.1 Motivating Example
Here, we provide an experimental analysis of the baselines’ exploration capacity on Navigation environment (cf. Fig.˜3), where the point-mass agent is controlled by an axis-aligned 222-dim velocity controller. We compare MTP-Bspline and MTP-Akima to evolutionary algorithms and standard MPC baselines, with maximum sampling noise settings (cf. Fig.˜3). In particular, given the control limits [1,1][-1,1][ - 1 , 1 ] on x-y axes, we set the standard deviation $\sigma=1$ for sampling noise of MPPI and PS, and population generation noise for OpenAI-ES and DE.
Figure 3: Motivation comparison of MTP methods versus baselines with $B=256$ on Navigation environment. The environment is designed to be challenging to reach the green goal, requiring strong exploration to avoid large local minima in the middle (see task details in Section˜A.3). We plan with $T=20$ with $\Delta t=0.05s$. The figures show 555 random traces of white rollouts. (Left) MTP-Akima rollouts reach the green goal very early due to high-entropy tensor sampling, while (Right) OpenAI-ES struggles to generate a rollout exploring the way out of large local minima.
Fig.˜3 also shows the cost convergence and the cost entropy curves over timesteps, in which the entropy is computed as $H=-\sum_{j=1}^{B}P_{j}\log P_{j},\,P_{j}=\exp(J_{j})/(\sum_{l}\exp(J_{l}))$, where $\{J_{j}\}_{j=1}^{B}$ is the batch of cumulative rollout costs. The entropy represents the diversity of rollout evaluation, implying the exploration capability of algorithms. We observe that MTP-Akima has the highest entropy curve, inducing the lowest cost convergence over timesteps, while other baselines converge to the middle minima. In this case, MTP-Bspline also struggles due to conservative interpolation in tensor sampling, achieving moderate entropy, yet higher than evolutionary strategies.
@@ 4.2 Comparison Experiments
We analyze the performance comparison of MTP and the baselines over various robotics tasks representing different dynamics and planning cost settings (cf. Fig.˜4). In each task, we tune the baselines and set the same white noise standard deviation for MTP/MPPI/PS to study the performance gain by tensor sampling, while using the default hyperparameters for evolutionary baselines.
Comparison Environments. PushT (Chi et al., 2023), Cube-In-Hand (Andrychowicz et al., 2020) require intricate manipulation environments where robust exploration is critical due to complex contact dynamics and precise multi-step manipulation requirements. G1-Standup, G1-Walk represent high-dimensional robotic tasks demanding substantial computational resources and sophisticated control strategies. Crane, Walker (Towers et al., 2024) present underactuated and nonlinear dynamic challenges. To ensure fair comparison, for all baselines, we fix the same number of rollouts $B=16$ on Crane, and $B=128$ for all other tasks. All tasks are implemented in hydrax (Kurtz, 2024). Further experiment details are in Section˜A.4.
Figure 4: Performance comparison of MTP variants against standard MPC methods (MPPI, PS) and evolutionary algorithms (OpenAI-ES, DE). The (horizontal) gray dashed line depicts the task success, while the (vertical) red line represents the timestep such that the first algorithm statistically succeeds the task, or fails last in G1-Walk.
@@
In the Crane environment, we apply heavy modeling errors of mass, inertia, and pulley/joint damping. MTP variants excel by maintaining stable control trajectories with smooth transitions, effectively navigating the nonlinear dynamics and underactuation. The B-spline interpolation’s conservative nature helps avoid overshooting and instability prevalent in these tasks, thus outperforming both evolutionary algorithms and standard MPC methods that tend to produce erratic control inputs. The Walker task exhibits a rather simple dynamics model and is less sensitive to the sampling distribution due to its relatively simple contact model. We apply no modeling error as a sanity check. Indeed, MTP and classical MPPI/PS perform similarly in simple cases.
@@ 4.3 Design Ablation
We conduct an ablation study analyzing the effect of varying two crucial hyperparameters—the number of elites $E$ and the mixing rate $\beta$—on MTP algorithmic performance. Fig.˜5 presents heatmaps indicating accumulated cost over time for each task and interpolation method (MTP-Linear, MTP-Bspline, MTP-Akima). Each heatmap illustrates distinct algorithmic realizations at its corners. Specifically, the bottom-left corner represents PS, characterized by a single elite and purely white noise sampling. Conversely, the bottom-right corner corresponds to Predictive Tensor Sampling (TS-PS), maintaining a single elite but employing full tensor sampling. Due to the softmax update (Algorithm˜2 Line 11-12), the top-left corner realizes MPPI (with adaptive sampling covariance), leveraging all candidate samples with local noise sampling, while the top-right corner reflects Tensor Sampling-MPPI (TS-MPPI), utilizing all candidates and full tensor sampling for maximum exploration.
Figure 5: Mixing rate $\beta$ and number of elites $E$ sweep on PushT, G1-Walk tasks with $B=128$ to investigate the algorithmic update rule Algorithm˜2 Line 9-12. The heatmap indicates accumulated cost over timesteps at termination, and the heat value range is fixed for each task/row.
We observe the consistent pattern that MTP performance degrades at the extremes of $\beta$. In PushT, moderate values of $\beta$ lead to significantly lower costs, while the absence of tensor sampling ($\beta=0$) yields poor performance due to inadequate exploration. This observation reinforces the effectiveness of the $\beta$-mixing strategy for balancing global and local sampling contributions. Interestingly, the number of elites $E$ has a limited impact in PushT, likely due to the task’s insensitivity to control stability. However, in G1-Walk, the choice of $E$ is crucial. Using a single elite ($E=1$), corresponding to the PS control scheme, leads to unstable and jerky behavior, which aligns with the poor performance observed in Fig.˜4. At the other extreme, with $\beta=1$ (full tensor sampling), performance also degrades. This is attributed to the fixed rollout budget $B$; as $M$ increases, global samples become too sparse to effectively capture the fine-grained control required for stable gait tracking. In this case, exploitation with local samples is essential to maintain intricate motion tracking control.
@@ 4.4 Sensitivity Ablation
Here, we perform sensitivity analyses for various critical algorithmic hyperparameters. Fig.˜6 evaluates sensitivity to the mixing rate $\beta$ for different tasks. For the PushT environment, results show minimal sensitivity across mixing rates, as the task inherently lacks significant failure modes, ensuring consistent success regardless of the exploration-exploitation balance. In contrast, the Cube-In-Hand task demonstrates high sensitivity, with larger mixing rates causing instability due to the cube falling out of grasp frequently. Optimal performance is thus achieved with lower $\beta$ values, suggesting careful management of exploration intensity. Furthermore, for the high-dimensional G1-Standup task, a smaller mixing rate helps stabilize the control, enabling the robot to achieve a more consistent and stable stand-up performance.
@@ 5 Related Works
We review related efforts across two major directions: the vectorization of sampling-based MPC and sampling-based motion planning. While these approaches originate from different planning paradigms, dynamics-aware MPC versus collision-free geometric planning, they share a common structure: interacting with an agent-environment model (e.g., dynamics model or collision checker) that can be vectorized for efficient, batched computation. Both domains benefit from high-throughput sampling, making them increasingly amenable to modern GPUs/TPUs.
@@
Sampling-based MPC Vectorization. Sampling-based MPC (Mayne, 2014) has been successfully applied to high-dimensional, contact-rich control problems. Methods such as Predictive Sampling (PS) (Howell et al., 2022), Model Predictive Path Integral (MPPI) (Williams et al., 2017; Watson & Peters, 2023), and CEM-based MPC (Pinneri et al., 2021) rely on parallel sampling of control trajectories and subsequent rollouts using a system dynamics model. These methods naturally benefit from vectorized simulation backends, and recent works have extended them toward more structured and efficient exploration. For instance, inspired by the diffusion process, DIAL-MPC (Xue et al., 2024) enhances exploration coverage and local refinement simultaneously, achieving high-precision quadruped locomotion and outperforming reinforcement learning policies (Schulman et al., 2017) in climbing tasks. STORM (Bhardwaj et al., 2022) demonstrates GPU-accelerated joint-space MPC for robotic manipulators, achieving real-time performance while handling task-space and joint-space constraints. Other recent efforts integrate GPU-parallelizable simulators, such as IsaacGym (Makoviychuk et al., 2021), into the MPC loop (Pezzato et al., 2025) for online domain randomization, removing the need for explicit modeling and enabling real-time contact-rich control. In another line, CoVO-MPC (Yi et al., 2024) enhances convergence speed by optimizing the covariance matrix during sampling, leading to performance gains in both simulated and real-world quadrotor tasks. These advances demonstrate that structured, parallel control sampling can be effectively deployed in high-stakes robotics applications using vectorized dynamic models.
@@
Motion Planning Vectorization. Recent advances in sampling-based motion planning have demonstrated that classical methods, such as RRT (Kuffner & LaValle, 2000), can be significantly accelerated through parallel computation, while preserving theoretical guarantees like probabilistic completeness, which represents another form of maximum exploration. Early work focused on accelerating specific subroutines like collision checking (Bialkowski et al., 2011; Pan & Manocha, 2012), but more recent efforts have restructured planners for full GPU-native execution. Examples include GMT* (Lawson et al., 2020), VAMP (Thomason et al., 2024), pRRTC (Huang et al., 2025), and Kino-PAX (Perrault et al., 2025), which achieve millisecond-scale planning in high-dimensional configuration spaces by parallelizing sampling, forward kinematics, and tree expansions. GTMP (Le et al., 2025) pushes this even further by implementing the sampling, graph-building, and search pipeline as tensor operations over batch-planning instances, showcasing the feasibility of real-time planning across multiple environments.
@@ 6 Discussions and Conclusions
In this work, we introduced Model Tensor Planning (MTP), a robust sampling-based MPC approach designed to achieve global exploration via maximum entropy sampling. Theoretically, we demonstrated that in the limits of infinite layers $M$ and samples per layer $N$, our tensor sampling method attains maximum entropy, thereby efficiently approximating the full trajectory space. Furthermore, MTP is intentionally designed to be practically feasible, enabling straightforward implementation for sampling high-entropy control trajectories (see Section˜A.5).
@@
Limitations
While MTP demonstrates strong performance across diverse control tasks, it inherits several limitations typical of sampling-based methods. First, its computational cost scales with the number of rollouts, making it challenging to deploy on hardware with limited parallel computing. Second, while our tensor-based sampler improves exploration coverage, it does not leverage task-specific priors or learning-based proposal distributions, which could further improve sample efficiency. Finally, MTP relies on a fixed dynamics model, limiting its robustness in partially observed or stochastic environments.
Broader Impact Statement
This work contributes to the development of efficient sampling-based control by introducing a scalable, high-entropy sampling mechanism for model predictive control (MPC). Model Tensor Planning (MTP) opens a promising direction in the design of exploratory algorithms that go beyond local refinements, allowing for global reasoning over control spaces. By enabling maximum entropy exploration via structured tensor operations, MTP provides a framework that may benefit a wide range of decision-making systems requiring robust performance in underexplored, high-dimensional environments, such as dexterous manipulation, legged locomotion, or autonomous vehicles operating under partial observability and uncertainty.
From a real-world deployment perspective, MTP maintains controls within physical limits, but its high-rate, tensorized control sequences may induce rapid variations that practical motor systems must robustly execute. While this fast-changing control is beneficial for agility and responsiveness, it necessitates attention to actuator dynamics and hardware safety. Therefore, safety-aware control filtering or actuator-aware smoothness constraints may be incorporated as extensions for deployment.
Furthermore, like other MPC approaches, MTP assumes access to reliable state estimation for initializing planning rollouts in the control loop. In practical deployment, this typically requires a real-time state estimator and a simulation back-end that serves as a digital twin. For example, MuJoCo XLA can simulate hundreds of dynamics instances in parallel, making it suitable for real-time predictive control. However, realizing this in hardware introduces engineering challenges, such as ensuring low-latency communication between the physical robot and the simulator. We see this digital twin architecture as a promising frontier where algorithmic advances like MTP can be tightly integrated with system-level design for robust, real-time, and scalable autonomous control.
@@ A.1 Proof of Theorem 1
Let $u\in{\mathbb{F}}$ be any path that is uniformly continuous and has bounded variation $TV(u)<\infty$. We begin by constructing a piecewise linear path approximating $u$, $g_{M}:[0,1]\to{\mathbb{U}}$ by dividing the interval [0,1][0,1][ 0 , 1 ] into $M-1$ subintervals, i.e., $[t_{1},t_{2}],\ldots,[t_{M-1},t_{M}]$ with $0=t_{1}<t_{2}<\ldots<t_{M}=1$. On each subinterval $[t_{i},t_{i+1}]$, we define the corresponding segment of $g$ to approximate $u$
@@
Since $u$ is uniformly continuous on [0,1][0,1][ 0 , 1 ], for any $\epsilon>0$, there exists $\delta>0$ such that for all $t,s\in[0,1]$, if $|t-s|<\delta$, then
@@
Let $u\in{\mathbb{F}}$ be any control path and ${\mathcal{G}}(M,N)$ be a random multipartite graph with $M$ layers and $N$ uniform samples per layer (cf. Definition˜1). Assuming a time sequence (i.e., knots) $0=t_{1}<t_{2}<\ldots<t_{M}=1$ with equal intervals, associating with layers $L_{1},\ldots,L_{M}\in{\mathcal{G}}(M,N)$ respectively, then
@@
First, Lemma˜2 implies that there exists a sequence of linear piecewise $g_{M}$, having $M-1$ equal intervals approximating $u$, converging to $u$ as $M\to\infty$.
Let $\hat{g}_{M}\in{\mathcal{G}}(M,N)$ be a control path in ${\mathcal{G}}$ (cf. Definition˜4). Since the time sequence $0=t_{1}<t_{2}<\ldots<t_{M}=1$ corresponding to layers $L_{1},\ldots,L_{M}\in{\mathcal{G}}(M,N)$ has equal intervals, we can consider $\hat{g}_{M}$ having $M-1$ segments approximating $g_{M}$ without loss of generality.
@@
for some $c>0$. From Lemma˜1, the probability that every layer contains at least one such sample, such that $\|g_{M}-\hat{g}_{M}\|_{\infty}<\epsilon$, is at least $1-Me^{-cN}$, which converges to 1 as $N\to\infty$.
From Lemma˜2, for sufficiently large $M$, we have $\|u-g_{M}\|_{\infty}<{\epsilon}$. Now, due to ${\mathbb{U}}$ is compact, we can apply the triangle inequality as $N\to\infty$